If P(A) = 0.7, P(B) = 0.8, and P(A or B) = 0.9, find P(A and B).
P(A and B) =_____

Answer by ikleyn(52747)

.

    P(A and B) = P(A) + P(B) - P(A or B).


It is one of the basic formulas in Probability theory.


Substitute the given data and calculate


    P(A and B) = 0.7 + 0.8 - 0.9 = 0.6.


Solved, answered and explained.